If the molecular weight of a molecule was 450 daltons, how much would be added to 100mL of water to make a 2 M solution?

If the molecular weight if a molecule is 400 daltons, how much would be added to 500 mL of water to make a 0.5 solution?
Answer #1

Formulas which we will use in this question :-

(1) Molarity (M) = Moles of solute/ Volume of solution (Litres)

(2) Moles = Given mass / Molar mass.

*Here given mass means, mass of solute which is used in formation of solution.

Now in question, First we will calculate number of moles and then we will get mass of solute.

Question 1.
Molarity = 2M
Volume = 100mL or 100/1000 = 0.1Litre
Molecular weight = 450 Daltons.

Using formula (1).

2 = Moles/ 0.1
So, Moles = 2×0.1 = 0.2

Now, Using formula (2).

0.2 = Mass of solute / 450.
, Mass of solute = 90 daltons.
So, 90 daltons of Solute/ Substance should be added.

Similarly for question 2.
Molarity = 0.5 M
Volume = 500 mL or 500/1000 = 0.5 Litres.
Molecular weight = 400 Daltons.

Using formula (1).

0.5 = Moles / 0.5
So, Moles = 0.5×0.5 = 0.25 Moles.

Now, Using formula (2).

0.25 = Mass of solute / 400
, Mass of solute = 0.25×400 = 100 Daltons.
So, 100 Daltons of Solute / Substance should be added.

* If you have to calculate answer in grams then
1 Dalton = 1.661 × 10^-24 gram